OpenStudy (anonymous):

A raffle ticket costs $10. 10 tickets net $500 and one ticket nets $2,000. The expected value of a ticket is $5. How many tickets were sold?

OpenStudy (anonymous):

\[\frac{10\times 500+2000-10\times n}{n}=5\] solve for \(n\)

OpenStudy (anonymous):

i hope it is clear how to solve the expected value is the amount you get times the probability you get it, added up i.e. the average payoff over all the tickets

OpenStudy (anonymous):

Thanks!